the question is find the solution to each system by the substitution method
2x - 5y = -1
x - 3y = -3

Answer by jorel1380(3719)

2x - 5y = -1
x - 3y = -3
x=3y-3
2(3y-3)-5y=-1
6y-6-5y=-1
y=-1+6=5
x=12..